# Offline Goal-Conditioned Reinforcement Learning with Projective Quasimetric Planning

### v2

@@ D.1.2 Discussion of Standing Assumptions
To the best of our knowledge, most navigation tasks in the literature, and all in OGBench [49], can be restricted to a bounded subset of $\mathbb{R}^{n}$. Enforcing that $\mathcal{S}$ is closed and bounded, and consequently compact in finite dimension, simply means there are well-defined "walls” or “obstacles” beyond which the agent never ventures. In the majority of simulated mazes, for example, the reachable region is a finite polygonal domain.
@@
We assume that our negative sampler can produce, for every point $s\in\mathbb{R}^{n}$ whose distance to $\mathcal{D}$ is at most $Diam(\mathcal{D})$, at least one generated sample $\tilde{s}$ within $\epsilon$. In other words, the set $\mathcal{\widetilde{D}}$ is $\epsilon$-dense in the controlled shell $\{\leavevmode\nobreak\ s\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ s\in\mathbb{R}^{n}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \min_{s^{\prime}\in\mathcal{D}}\leavevmode\nobreak\ \lVert s-s^{\prime}\rVert_{2}\leq Diam(\mathcal{D})\leavevmode\nobreak\ \}$. Concretely, one can implement this by drawing uniform random points in the ball of radius $Diam(\mathcal{D})$ around each $s\in\mathcal{D}$. In low dimension ($n=2$ or $n=3$), one can use a grid to guarantee $\epsilon$-covering.
@@
MLPs with Lipschitz activations (such as ReLU, GELU, or even Sigmoid), and eventually layer normalization, are provably Lipschitz [65]. For example in a MLP with K hidden layers, if every weight matrix $W_{i}$ has $\lVert W_{i}\rVert\leq\kappa$ and each activation is $1$-Lipschitz, then the MLP itself is L-Lipschitz, with $L=\kappa^{K}$. In practice, one can enforce weight clipping, weight regularization, spectral-norm regularization, or to keep the Lipschitz coefficient relatively small [43].
@@
During training of the classifier $\psi$, each positive state $s\in\mathcal{D}$ incurs a loss $-\lambda_{\text{ood}}\cdot\log\big{(}\psi(\phi(s))\big{)}$ with $\lambda_{\text{ood}}$ steadily increased until $\psi(\phi(s))\geq 1-\delta$ for all $s$. That is a simple, yet effective, Lagrangian strategy to force $\psi$ into $[1-\delta,1]$ for positive samples.
@@
By construction, any $\tilde{s}\in\mathcal{\widetilde{D}}$ incurs loss $-\log\big{(}1-\psi(\phi(\tilde{s}))\big{)}$ which pushes $\psi(\phi(\tilde{s}))$ toward 00. We define $\bar{\delta}=\max\{\leavevmode\nobreak\ \psi(\phi(\tilde{s}))\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \tilde{s}\in\mathcal{\widetilde{D}}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \eta\leq\min_{s\in\mathcal{D}}\leavevmode\nobreak\ \lVert\tilde{s}-s\rVert_{2}\}$.In practice, one cannot guarantee $\bar{\delta}=0$ exactly, but empirical training drives $\bar{\delta}$ to a small value.
@@ D.1.3 Discussion of Limitations
High-Dimensional State Spaces (large n) :
In $\mathbb{R}^{n}$ with $n>2$, uniformly covering the previously defined shell up to a fine $\epsilon$ becomes exponentially costly in $n$. Grid-based sampling will require $O(\epsilon^{-n})$ points. In high-dimensional pixel-based observations (e.g. $n\simeq 1000$), one cannot hope to sample negatives points in that entire ball. Instead we should exploit structures such as known game map geometry, or pre-trained world models to propose out-of-distribution samples more intelligently.
@@
Our proposed ProQ implementation use one-step linear interpolation $(1-\alpha)\cdot s_{1}+\alpha\cdot s_{2}$ and extrapolation $(1+\beta)\cdot s_{1}-\beta\cdot s_{2}$, with $\alpha,\beta\sim\mathcal{U}([0,1])$ and $s_{1},s_{2}\in\mathcal{D}$. These do not uniformly cover the entire negative shell in $\mathbb{R}^{n}$ for $n>2$, they only explore the convex hull of pairs and its immediate outward extensions. Hence, there may exist unreachable points $s$ within the admissible shell that never get generated, leaving $\psi$ unconstrained. A uniform-shell sampler would ensure no such gap, but is expensive when $n$ is large.
@@ D.2 Coulomb-like Energy & Uniform Keypoint Spreading
In this appendix, we discuss and analyze why minimizing a Coulomb-like repulsive energy over a finite set of learnable keypoints $\{p_{i}\}_{i=1}^{K}\subset\mathbb{R}^{d}$ pushes those keypoints to spread uniformly over a compact domain $\phi(\mathcal{S})$. We begin by recalling the classical Coulomb law in physics, then show how our loss in ProQ reduces to a pure repulsive energy in latent space. Finally, we prove under mild assumptions that any minimizer of this energy must maximize pairwise separation and, asymptotically, approaches an uniform distribution.
@@ D.2.1 From Coulomb’s Law to Our Energy Loss
In electrostatics, two positively-charged particles located at $x_{i},x_{j}\in\mathbb{R}^{3}$ experience a potential :
$V(x_{i},x_{j})=\frac{q_{i}q_{j}}{4\pi\epsilon_{0}}\cdot\frac{1}{\lVert x_{i}-x_{j}\rVert_{2}}\leavevmode\nobreak$ (17)
@@
$E(x_{1},\ldots,x_{K})=\sum_{1\leq i<j\leq K}\frac{1}{\lVert x_{i}-x_{j}\rVert_{2}}\leavevmode\nobreak\ .$ (18)
Minimizing this energy $E(x_{1},\ldots,x_{K})$ on a compact domain (e.g. the unit sphere $\mathbb{S}^{2}$) yields the classical Thomson Problem [7, 33] of finding well-separated, nearly-uniform point configurations.
In ProQ, we replace Euclidean distance by a learned latent quasimetric $d:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$. Concretely, if $\{z_{k}\}_{k=1}^{K}$ are the embeddings of the latent learnable keypoints, then our Coulomb-like repulsive loss becomes :
$\mathcal{L}_{\mathrm{repel}}=\sum_{i\neq j}\frac{1}{d(z_{i},z_{j})+\epsilon}\leavevmode\nobreak\ ,$ (19)
where $\epsilon>0$ is a small constant that prevents numerical blow-up when two points become very close.
@@
with $\lambda_{\text{ood}}$ updated until satisfaction of a penalty $\big{(}(1-\delta)-\sum_{k=1}^{K}\psi(z_{k})\big{)}^{2}$, with $10<\delta<$ 1 the compliance. In other words, on average, all latent keypoints must stay in-distribution to some extent.
@@ D.2.2 Theoretical Analysis
    $\exists\leavevmode\nobreak\ c_{1},c_{2}>0,\leavevmode\nobreak\ \forall x,y\in\mathcal{S},\leavevmode\nobreak\ c_{1}\cdot\lVert x-y\rVert_{2}\leq d(x,y)\leq c_{2}\cdot\lVert x-y\rVert_{2}$ (21)
@@
    $E:\mathcal{S}^{K}\rightarrow\mathbb{R}^{+}\cup\{+\infty\}\leavevmode\nobreak\ ,\ s_{1},\ldots,s_{K}\rightarrow\sum_{i\neq j}\frac{1}{d(s_{i},s_{j})}$ (22)
  • A metric $d^{\prime}:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}^{+}$ defined as : $\forall x,y\in\mathcal{S},\leavevmode\nobreak\ d^{\prime}(x,y)=\min\big{(}d(x,y),d(y,x)\big{)}$
@@
Existence and Distinctness of Energy Minimizers : Let’s consider the energy function $E:\mathcal{S}^{K}\rightarrow\mathbb{R}^{+}\cup\{+\infty\}$ Under the previous assumptions of compactness and continuity we have :
@@
(ii) For any such minimizer $\{s^{*}_{k}\}_{k=1}^{K}\subset\mathcal{S}\leavevmode\nobreak$, if $K>1$ then all keypoints are distinct.
@@
If $\mathcal{S}\subset\mathbb{R}^{n}$ is reduced to a single point, the proof is direct (the energy is 00 as the empty sum).
If $\mathcal{S}$ is not reduced to a single point, we can select a configuration $\{u_{k}\}_{k=1}^{K}\subset\mathcal{S}$ of $K$ distinct points. Then $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ u_{i}\neq u_{j}$, we have $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(u_{i},u_{j})\neq 0$ and $M_{o}=E(u_{1},\ldots,u_{K})>0$ is finite.
We consider the set $\mathcal{C}=\big{\{}\{s_{k}\}_{k=1}^{K}\in\mathcal{S}^{K}\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ E(s_{1},\ldots,s_{K})\leq M_{0}\big{\}}$.
@@
$\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ \frac{1}{d(s_{i},s_{j})}\leq\sum_{i^{\prime}\neq j^{\prime}}\frac{1}{d(s_{i^{\prime}},s_{j^{\prime}})}\leq M_{0}$ (24)
@@
$\forall\leavevmode\nobreak\ \{s_{k}\}_{k=1}^{K}\in\mathcal{C}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(s_{i},s_{j})\geq\frac{1}{M_{0}}$ (25)
Let’s consider ($\big{(}\{s_{k}^{(n)}\}_{k=1}^{K}\big{)}_{n\in\mathbb{N}}$ be a converging sequence in $\mathcal{C}$ with a limit $\{s_{k}^{*}\}_{k=1}^{K}\in\mathcal{S}^{K}$.
@@
$\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(s_{i}^{(n)},s_{j}^{(n)})\geq\frac{1}{M_{0}}\Rightarrow\lim_{n\rightarrow+\infty}d(s_{i}^{(n)},s_{j}^{(n)})=d(s_{i}^{*},s_{j}^{*})\geq\frac{1}{M_{0}}$
@@
Lemma D.5.
Existence of a Well-Separated Configuration : Considering the compact metric space $(\mathcal{S},d^{\prime})$, there exists … such that for any integer $K\in\mathbb{N}^{*}$ we can find a configuration $\{s_{k}\}_{k=1}^{K}\subset\mathcal{S}$ verifying : $\min_{1\leq i<j\leq K}\leavevmode\nobreak\ d^{\prime}(s_{i},s_{j})>\delta_{K}\leavevmode\nobreak\ ,$ where $\delta_{K}=\leavevmode\nobreak\ ???$
Proof.
Let $\delta>0$ be a separation distance. We construct a finite set $U=\{u_{k}\}_{k=1}^{M}\subset\mathcal{S}$ as it follows :
  1. 1.
    Initialization : Pick an arbitrary point $u_{1}\in\mathcal{S}$ and set $U=\{u_{1}\}$ and $M=1$ ;
  2. 2.
    Iteration : While $\exists\leavevmode\nobreak\ u\in\mathcal{S}$ such that $\forall\leavevmode\nobreak\ u_{k}\in U\leavevmode\nobreak\ ,\leavevmode\nobreak\ d^{\prime}(u,u_{k})\geq\delta\leavevmode\nobreak$, set $u_{M+1}=u$ and increment $M$.
  3. 3.
    Termination : If no such point $u$ can be find, the final set $U=\{u_{k}\}_{k=1}^{M}$ is obtained.
The algorithm terminates, otherwise we can generate an infinite sequence =$=\{u_{n}\}_{n\in\mathbb{N}}\subset\mathcal{S}$ with $d^{\prime}(u_{i},u_{j})\geq\delta>0$ for all $i\neq j$. However, since $\mathcal{S}$ is compact, this sequence would have a convergent subsequence which lead to $\exists\leavevmode\nobreak\ i,j$ such that $d^{\prime}(u_{i},u_{j})<\delta$, contradicting our construction process. Thus $M$ is finite, $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d^{\prime}(u_{i},u_{j})\geq\delta$. Moreover the set $U\subset\mathcal{S}$ is maximally separated and covers $\mathcal{S}$, because if not the case we could have extended $U$ by adding $u\in\mathcal{S}$ further from at least $\delta$.
@@ D.3 Learned State-Space Mappings
Figure 8 shows the learned state mappings. In the Medium maze, corridors are consistently identified as in-distribution, and the red keypoints uniformly occupy each hallway. We notably notice that varying the seed does not significantly alter the learned latent coverage. Similarly, in the Large maze, the OOD boundary tightly traces the walls, and keypoints remain evenly spaced along identical corridors for all seeds, confirming robustness. In the Teleport environment, the portal squares introduce stochastic transitions ; the OOD classifier occasionally colors portal cells green (intermediate confidence), and keypoints cluster variably near these portals for different seeds, indicating sensitivity to random initialization in regions with non-deterministic dynamics. Finally, in the Giant maze, the expansive corridor network is consistently recognized as in-distribution, and keypoints again spread uniformly across all three seeds, illustrating that ProQ’s mapping remains stable even at larger scale.
@@
Figure 8: Learned State-Space Mappings. Each subplot displays the OOD classifier’s output (from yellow for high OOD probability, to purple for high in-distribution probability) along with the learned keypoints as red dots. Each of the rows correspond to a different environment (Medium, Large, Teleport, and Giant), and each of the columns correspond to a different seeding ($100$, $200$, and $300$).

### v3

@@ D.1.2 Discussion of Standing Assumptions
To the best of our knowledge, most navigation tasks in the literature, and all in OGBench [49], can be restricted to a bounded subset of $\mathbb{R}^{n}$. Enforcing that $\mathcal{S}$ is closed and bounded, and consequently compact in finite dimension, simply means there are well‐defined "walls” or “obstacles” beyond which the agent never ventures. In the majority of simulated mazes, for example, the reachable region is a finite polygonal domain.
@@
We assume that our negative sampler can produce, for every point $s\in\mathbb{R}^{n}$ whose distance to $\mathcal{D}$ is at most $Diam(\mathcal{D})$, at least one generated sample $\tilde{s}$ within $\epsilon$. In other words, the set $\mathcal{\widetilde{D}}$ is $\epsilon$-dense in the controlled shell $\{\penalty 10000\ s\penalty 10000\ \big|\penalty 10000\ s\in\mathbb{R}^{n}\penalty 10000\ ,\penalty 10000\ \min_{s^{\prime}\in\mathcal{D}}\penalty 10000\ \lVert s-s^{\prime}\rVert_{2}\leq Diam(\mathcal{D})\penalty 10000\ \}$. Concretely, one can implement this by drawing uniform random points in the ball of radius $Diam(\mathcal{D})$ around each $s\in\mathcal{D}$. In low dimension ($n=2$ or $n=3$), one can use a grid to guarantee $\epsilon$‐covering.
@@
MLPs with Lipschitz activations (such as ReLU, GELU, or even Sigmoid), and eventually layer normalization, are provably Lipschitz [65]. For example in a MLP with K hidden layers, if every weight matrix $W_{i}$ has $\lVert W_{i}\rVert\leq\kappa$ and each activation is $1$-Lipschitz, then the MLP itself is L‐Lipschitz, with $L=\kappa^{K}$. In practice, one can enforce weight clipping, weight regularization, spectral‐norm regularization, or to keep the Lipschitz coefficient relatively small [43].
@@
During training of the classifier $\psi$, each positive state $s\in\mathcal{D}$ incurs a loss $\lambda_{\text{ood}}\cdot\log\big(\psi(\phi(s))\big)$ with $\lambda_{\text{ood}}$ steadily increased until $\psi(\phi(s))\geq 1-\delta$ for all $s$. That is a simple, yet effective, Lagrangian strategy to force $\psi$ into $1-\delta,1]$ for positive samples.
@@
[…]
@@ D.1.3 Discussion of Limitations
High‐Dimensional State Spaces (large n) :
In $\mathbb{R}^{n}$ with $n>2$, uniformly covering the previously defined shell up to a fine $\epsilon$ becomes exponentially costly in $n$. Grid‐based sampling will require $O(\epsilon^{-n})$ points. In high‐dimensional pixel‐based observations (e.g. $n\simeq 1000$), one cannot hope to sample negatives points in that entire ball. Instead we should exploit structures such as known game map geometry, or pre‐trained world models to propose out‐of‐distribution samples more intelligently.
@@
Our proposed ProQ implementation use one‐step linear interpolation $(1-\alpha)\cdot s_{1}+\alpha\cdot s_{2}$ and extrapolation $(1+\beta)\cdot s_{1}-\beta\cdot s_{2}$, with $\alpha,\beta\sim\mathcal{U}([0,1])$ and $s_{1},s_{2}\in\mathcal{D}$. These do not uniformly cover the entire negative shell in $\mathbb{R}^{n}$ for $n>2$, they only explore the convex hull of pairs and its immediate outward extensions. Hence, there may exist unreachable points $s$ within the admissible shell that never get generated, leaving $\psi$ unconstrained. A uniform‐shell sampler would ensure no such gap, but is expensive when $n$ is large.
@@ D.2 Coulomb-like Energy & Uniform Keypoint Spreading
In this appendix, we discuss and analyze why minimizing a Coulomb‐like repulsive energy over a finite set of learnable keypoints $\{p_{i}\}_{i=1}^{K}\subset\mathbb{R}^{d}$ pushes those keypoints to spread uniformly over a compact domain $\phi(\mathcal{S})$. We begin by recalling the classical Coulomb law in physics, then show how our loss in ProQ reduces to a pure repulsive energy in latent space. Finally, we prove under mild assumptions that any minimizer of this energy must maximize pairwise separation and, asymptotically, approaches an uniform distribution.
@@ D.2.1 From Coulomb’s Law to Our Energy Loss
In electrostatics, two positively‐charged particles located at $x_{i},x_{j}\in\mathbb{R}^{3}$ experience a potential :
$V(x_{i},x_{j})=\frac{q_{i}q_{j}}{4\pi\epsilon_{0}}\cdot\frac{1}{\lVert x_{i}-x_{j}\rVert_{2}}\penalty 10000\$(17)
@@
$E(x_{1},\ldots,x_{K})=\sum_{1\leq i<j\leq K}\frac{1}{\lVert x_{i}-x_{j}\rVert_{2}}\penalty 10000\ .$ (18)
Minimizing this energy $E(x_{1},\ldots,x_{K})$ on a compact domain (e.g. the unit sphere $\mathbb{S}^{2}$) yields the classical Thomson Problem [7, 33] of finding well‐separated, nearly‐uniform point configurations.
In ProQ, we replace Euclidean distance by a learned latent quasimetric $d:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$. Concretely, if $\{z_{k}\}_{k=1}^{K}$ are the embeddings of the latent learnable keypoints, then our Coulomb‐like repulsive loss becomes :
$\mathcal{L}_{\mathrm{repel}}=\sum_{i\neq j}\frac{1}{d(z_{i},z_{j})+\epsilon}\penalty 10000\ ,$ (19)
where $\epsilon>0$ is a small constant that prevents numerical blow‐up when two points become very close.
@@
with $\lambda_{\text{ood}}$ updated until satisfaction of a penalty $\big((1-\delta)-\sum_{k=1}^{K}\psi(z_{k})\big)^{2}$, with $0<\delta<1$ the compliance. In other words, on average, all latent keypoints must stay in-distribution to some extent.
@@ D.2.2 Theoretical Analysis
    $\exists\penalty 10000\ c_{1},c_{2}>0,\penalty 10000\ \forall x,y\in\mathcal{S},\penalty 10000\ c_{1}\cdot\lVert x-y\rVert_{2}\leq d(x,y)\leq c_{2}\cdot\lVert x-y\rVert_{2}$ (21)
@@
    $E:\mathcal{S}^{K}\rightarrow\mathbb{R}^{+}\cup\{+\infty\}\penalty 10000\ ,\ s_{1},\ldots,s_{K}\rightarrow\sum_{i\neq j}\frac{1}{d(s_{i},s_{j})}$ (22)
  • A metric $d^{\prime}:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}^{+}$ defined as : $\forall x,y\in\mathcal{S},\penalty 10000\ d^{\prime}(x,y)=\min\big(d(x,y),d(y,x)\big)$ .
@@
Existence and Distinctness of Energy Minimizers : Let’s consider the energy function $E:\mathcal{S}^{K}\rightarrow\mathbb{R}^{+}\cup\{+\infty\}$ . Under the previous assumptions of compactness and continuity we have :
@@
(ii) For any such minimizer $\{s^{*}_{k}\}_{k=1}^{K}\subset\mathcal{S}\penalty 10000\$, if $K>1$ then all keypoints are distinct.
@@
If $\mathcal{S}\subset\mathbb{R}^{n}$ is reduced to a single point, the proof is direct (the energy is 0 as the empty sum).
If $\mathcal{S}$ is not reduced to a single point, we can select a configuration $\{u_{k}\}_{k=1}^{K}\subset\mathcal{S}$ of $K$ distinct points. Then $\forall\penalty 10000\ i\neq j\penalty 10000\ ,\penalty 10000\ u_{i}\neq u_{j}$, we have $\forall\penalty 10000\ i\neq j\penalty 10000\ ,\penalty 10000\ d(u_{i},u_{j})\neq 0$ and $M_{o}=E(u_{1},\ldots,u_{K})>0$ is finite.
We consider the set $\mathcal{C}=\big\{\{s_{k}\}_{k=1}^{K}\in\mathcal{S}^{K}\penalty 10000\ \big|\penalty 10000\ E(s_{1},\ldots,s_{K})\leq M_{0}\big\}$.
@@
$\forall\penalty 10000\ i\neq j\penalty 10000\ ,\penalty 10000\ \frac{1}{d(s_{i},s_{j})}\leq\sum_{i^{\prime}\neq j^{\prime}}\frac{1}{d(s_{i^{\prime}},s_{j^{\prime}})}\leq M_{0}$ (24)
@@
$\forall\penalty 10000\ \{s_{k}\}_{k=1}^{K}\in\mathcal{C}\penalty 10000\ ,\penalty 10000\ \forall\penalty 10000\ i\neq j\penalty 10000\ ,\penalty 10000\ d(s_{i},s_{j})\geq\frac{1}{M_{0}}$ (25)
Let’s consider $\big(\{s_{k}^{(n)}\}_{k=1}^{K}\big)_{n\in\mathbb{N}}$ be a converging sequence in $\mathcal{C}$ with a limit $\{s_{k}^{*}\}_{k=1}^{K}\in\mathcal{S}^{K}$.
@@
$\forall\penalty 10000\ i\neq j\penalty 10000\ ,\penalty 10000\ d(s_{i}^{(n)},s_{j}^{(n)})\geq\frac{1}{M_{0}}\Rightarrow\lim_{n\rightarrow+\infty}d(s_{i}^{(n)},s_{j}^{(n)})=d(s_{i}^{*},s_{j}^{*})\geq\frac{1}{M_{0}}$
@@ D.3 Learned State-Space Mappings
Figure 8 shows the learned state mappings. In the Medium maze, corridors are consistently identified as in‐distribution, and the red keypoints uniformly occupy each hallway. We notably notice that varying the seed does not significantly alter the learned latent coverage. Similarly, in the Large maze, the OOD boundary tightly traces the walls, and keypoints remain evenly spaced along identical corridors for all seeds, confirming robustness. In the Teleport environment, the portal squares introduce stochastic transitions ; the OOD classifier occasionally colors portal cells green (intermediate confidence), and keypoints cluster variably near these portals for different seeds, indicating sensitivity to random initialization in regions with non‐deterministic dynamics. Finally, in the Giant maze, the expansive corridor network is consistently recognized as in‐distribution, and keypoints again spread uniformly across all three seeds, illustrating that ProQ’s mapping remains stable even at larger scale.
@@
Figure 8: Learned State-Space Mappings. Each subplot displays the OOD classifier’s output (from yellow for high OOD probability, to purple for high in‐distribution probability) along with the learned keypoints as red dots. Each of the rows correspond to a different environment (Medium, Large, Teleport, and Giant), and each of the columns correspond to a different seeding ($100$, $200$, and $300$).